OpenStudy (anonymous):

Help please!!! A technician determined the thickness of zinc on a galvanized washer. He collected these data. outer diameter of washer, mm =26.4 inner diameter of washer, mm =10.8 height of washer, mm=2.2 mass of washer before reaction, g= 8.473 mass of washer after reaction, g=7.957 Questions: Calculate the volume in cubic mm of the zinc coating? Calculate the total surface area, in square millimeters (mm^2), of the washer Calculate the thickness, in millimeters, of the zinc coating.

OpenStudy (anonymous):

To start with you need to list out your givens and formulas, the process the often become rather clear and just Algebra where you're solving for unknowns and making substitutions. Treat the washer as the difference (subtraction of larger minus smaller) of the two volumes of cylinders V = V\(\large_{inner}\)-V\(\large_{outer}\) Volume of cylinder = 2\(\pi\)(radius)\(^2\)(height) Change in mass (aka. mass loss) = 8.473g - 7.957g Surface area of a cylinder = 2\(\pi\)(radius)\(^2\) + 2\(\pi\)(radius)(height) To find the zinc coating, you need to compare the change in mass to that of the density before and the density after. The density of zinc is known. \(\large7.14\frac{g}{cm^3}\)